# Primordial black hole capture, gravitational wave beats, and the nuclear equation of state

### v1

@@ Acknowledgements.
We would like to thank Steve Naculich and David Kaiser for helpful conversations. T.W.B. gratefully acknowledges the hospitality of the University of Arizona’s Steward Observatory. This work was supported in part by National Science Foundation (NSF) grants PHY-2010394 and PHY-2341984 to Bowdoin College, as well as NSF grants PHY-2006066 and PHY-2308242 to the University of Illinois at Urbana-Champaign.
@@
In this supplemental material we provide additional information on the construction of the stellar background models (Section I), the equations of motion (Section II), as well as the perturbation of nearly circular orbits (Section III). We adopt geometric units with $c=1=G$ throughout.
I Stellar background models
I.1 Oppenheimer-Volkoff solutions
@@
where $R$ is the areal radius. Here the metric coefficient $e^{2\lambda}$ is given by
@@
I.2 Transformation to isotropic coordinates
@@
In the stellar interior, however, Eq. (25) has to be integrated numerically in general, starting at the center with $r=0=R$ and fixing a constant of integration by matching to the exterior solution (I.2) at the stellar surface. We note that the singularity at $R=0$ on the right-hand side of (25) can be handled by adding and subtracting a term $dR/R$, i.e. by writing
@@
I.3 Constant-density star
Constant density stars are particularly useful since they can be constructed analytically and yield results for stars at the limit of stiffness, with $\Gamma=\infty$. For a constant-density star with $\rho=$ const we have
@@ I.3 Constant-density star
where $C=e^{\tilde{C}}$. We can now fix this constant by requiring that the isotropic radius in the interior, given by (I.3), matches its value in the exterior, given by (I.2), at the stellar surface $R_{*}$. This yields
@@
where $r_{*}=r(R_{*})$ is the isotropic stellar radius. We can also invert (I.3) to obtain
@@
II Primordial black holes as test particles
II.1 Geodesic equation
We track the primordial black hole’s (PBH’s) trajectory by treating it as a freely-falling test particle in the spacetime generated by a stellar background model as discussed in Section I. These spacetimes posses two Killing vectors – a time-like Killing vector $\xi_{(t)}=\partial/\partial t$ and a space-like Killing vector $\xi_{(\varphi)}=\partial/\partial\phi$ – which give rise to two constants of motion, namely the energy per unit mass
@@ II.1 Geodesic equation
Here $u^{a}$ is the PBH’s four-velocity, and the metric components can either be evaluated in terms of the areal radius $R$ (using the metric 16) or in terms of the isotropic radius $r$ (using the metric 23).
@@
(see, e.g., [29, 30] and note that the last equation is a direct consequence of 39). In (II.1), the time component of the four-velocity $u^{t}$ can either be found from (38), knowing $e$, or from the normalization $u_{a}u^{a}=-1$, which yields
@@
II.2 Initial conditions
We start all orbits by placing the PBH at an areal radius $R(0)=R_{\rm frac}R_{*}$ and $\varphi(0)=0$. From $R(0)$ we compute the initial isotropic radius $r(0)$ as described in Section I.2.
@@ II.2 Initial conditions
III Perturbations of circular orbits
@@ III.1 Stellar exterior
In this case several terms in (III) cancel conveniently and we obtain the more familiar expression,
@@
III.2 Stellar interior
In the stellar interior the analysis is more complicated because the metric coefficients $g_{tt}$ and $g_{RR}$ are no longer inverses of each other and hence no longer cancel each other. One approach would be to insert numerical solutions $M(R)$ and $\Phi(R)$ of the OV equations (see Section I) into (III). Instead, we focus our analysis to a nearly homogeneous region close to the center, where we expand the density and pressure about their central values [^2] The $R^{4}$ term in the expansion of $\rho(R)$ is needed in the $M(R)\ell^{2}/R^{3}$ term of (III) only.
@@ III.2 Stellar interior
where $\Phi_{0}=\Phi(0)$ is a constant of integration that is fixed by matching to the exterior solution at the stellar surface. We now insert $M(R)$ and $\Phi(R)$ into (III) and expand to order $R^{4}$ about the origin in order to identify the coefficients $p$ and $q$ in
@@
We next evaluate the derivatives $\rho_{c}^{(2)}$, $\rho_{c}^{(4)}$, and $P_{c}^{(2)}$ that we introduced in the expansions (III.2). In order to obtain $P_{c}^{(2)}$ we take a derivative of the stellar structure equation (20), replace the derivative of $M(R)$ using (19), and evaluate the result at $R=0$ to obtain
@@
As a consistency check we can take the Newtonian limit of (III.2) to find
@@
where we have used Newtonian expressions for the speed of sound $a^{2}=\Gamma K\rho^{\Gamma-1}$ as well as the Newtonian limit of (58), and where we do not distinguished between $\rho$ and $\rho_{0}$. As expected, (III.2) agrees with the expressions that we had found from Newtonian arguments in the main body of the paper. While it is instructive to consider the Newtonian limit, we note that the relativistic terms dominate the periastron advance for the neutron star hosts that we consider here, [^3] This can be verified by comparing the Newtonian terms with first-order post-Newtonian terms. so that a Newtonian treatment of these effects would lead to significant errors.
@@
by inserting the general relativistic expressions (60b) and (60c). Evaluating these terms requires the values ${\rho}_{c}$, $\rho_{0c}$, $P_{c}$, and $\Phi_{0}$, which depend on the stellar background model, as well as $E$ and $\ell$, which we determine for any given orbit. Results for specific stellar models and orbits are provided Table I of the main paper, where they are compared with results from integrations of the geodesic equations of Section II.
III.3 Constant-density star
For a constant-density star, both the mass $M(R)$ and the metric coefficient $e^{\Phi}$ are known analytically, see Eqs. (29) and (31). Accordingly, we can insert these expressions into (III) to obtain the potential $V(R)$ for all radii, without the need of an expansion about the origin. Abbreviating
@@ III.3 Constant-density star
for all radii. Unlike the purely Newtonian prediction, this fully relativistic calculation leads to a non-vanishing periastron advance even for constant-density stars, in agreement with our numerical solutions to the geodesic equation of Section II. For the orbits listed in Table I of the main paper, for example, for which $R_{\rm frac}=0.05$, $\ell_{\rm frac}=0.99$ and $M_{*}/R_{*}=1/6$, Eq. (65) yields a periastron advance of $\Delta\phi=0.001659$.
While, for a constant-density star, an expansion about $R=0$ is not necessary, such an expansion is instructive for purposes of comparison with the approach of Section III.2. In this case we obtain

### v2

@@ Acknowledgements.
We would like to thank Steve Naculich, David Kaiser, and Keith Riles for helpful conversations. T.W.B. gratefully acknowledges the hospitality of the University of Arizona’s Steward Observatory. This work was supported in part by National Science Foundation (NSF) grants PHY-2010394 and PHY-2341984 to Bowdoin College, as well as NSF grants PHY-2006066 and PHY-2308242 to the University of Illinois at Urbana-Champaign.
@@
In this supplemental material we provide additional information on expected event rates (Section I), the construction of the stellar background models (Section II), the equations of motion (Section III), as well as the perturbation of nearly circular orbits (Section IV).
I Event rates
Rates of collisions between neutron stars and PBHs have been estimated by a number of different authors (e.g. [30, 24, 25, 31, 32] and references therein). Here we briefly review some of the key arguments required for a rough estimate.
An individual neutron star will be hit by PBHs at a rate of approximately
$\dot{\mathcal{N}}_{\rm NS}\simeq\sigma v_{\infty}n_{\rm PBH},$ (16)
where $\sigma$ is the cross-section for the collision, $v_{\infty}$ the relative speed at large distances, and $n_{\rm PBH}$ the number density of the PBHs.
We estimate the cross-section $\sigma$ from
$\sigma=\pi R_{*}^{2}\left(1+\left(\frac{v_{\rm esc}}{v_{\infty}}\right)^{2}\right)$ (17)
(see, e.g., [30, 31]), where $v_{\rm esc}=(2GM_{*}/R_{*})^{1/2}$ is the escape speed from the neutron star with mass $M_{*}$ and radius $R_{*}$. For a typical neutron star with $M_{*}\simeq 1.4M_{\odot}$ and $R_{*}\simeq 6GM_{*}/c^{2}\simeq 12.5$ km we have $v_{\rm esc}\simeq 0.6\,c$. Further adopting $v_{\infty}\simeq 220$ km/s as a typical speed for halo dark-matter particles (see, e.g., [50]) we obtain
$\sigma\simeq 3\times 10^{18}\,\mbox{cm}^{2}.$ (18)
We take this value as a constant in the following, ignoring possible changes in $v_{\infty}$.
The PBH number density $n_{\rm PBH}$ is
$n_{\rm PBH}\simeq f_{\rm PBH}\frac{\rho_{\rm DM}}{m},$ (19)
where $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$ is the mass fraction of the Universe’s dark matter content in PBHs, $\rho_{\rm DM}$ is the mass density of dark matter, and we assume that all PBHs have the same mass $m$. Locally, in the solar neighborhood, the mass density of dark matter is approximately
$\rho_{\rm DM}^{\rm loc}\simeq 1.3\times 10^{-2}M_{\odot}\,\mbox{pc}^{-3}\simeq 10^{-24}\mbox{g}\,\mbox{cm}^{-3}$ (20)
(see, e.g., [51]).
Inserting the above expressions and numbers into (16) we obtain
$\dot{\mathcal{N}}_{\rm NS}\simeq 3\times 10^{-26}\,\mbox{s}^{-1}\,f_{\rm PBH}\left(\frac{10^{-6}M_{\odot}}{m}\right)\left(\frac{\rho_{\rm DM}}{\rho_{\rm DM}^{\rm loc}}\right).$ (21)
The Galactic collision rate is then approximately
$\dot{\mathcal{N}}_{\rm Gal}\simeq 10^{-9}\,\mbox{yr}^{-1}\,f_{\rm PBH}\left(\frac{10^{-6}M_{\odot}}{m}\right)\left(\frac{\rho_{\rm DM}^{\rm ave}}{\rho_{\rm DM}^{\rm loc}}\right)\left(\frac{N_{\rm NS}}{10^{9}}\right),$ (22)
where $N_{\rm NS}$ is the number of neutron stars in our Galaxy, and where $\rho_{\rm DM}^{\rm ave}$ represents a Galactic average of the dark-matter mass density. Near the Galactic center, in particular, the dark matter density is expected to be larger than its local value in the solar neighborhood by several orders of magnitude (see, e.g., [52]).
We note that the above event rates are inversely proportional to the PBH mass $m$, meaning that a smaller value of $m$ will result in a larger Galactic event rate. However, the strain $h$ of the emitted gravitational wave signal is proportional to $m$, so that the distance $d_{\rm max}$ to the furthest observable event, for a given detector sensitivity $h_{\rm min}$, is also proportional to $m$. Accordingly, the observable volume scales with $m^{3}$, and the total number of observable events – assuming a uniform distribution of potential sources – is proportional to $m^{2}$.
As a specific example, consider the orbits shown in Fig. 1 of the main paper, for which
$h$ $\simeq 10^{-24}\left(\frac{m}{10^{-6}M_{\odot}}\right)\left(\frac{10\,\mbox{kpc}}{d}\right).$ (23)
Given a detector sensitivity $h_{\rm min}$ for a continuous gravitational wave signal at a frequency of a few kHz we can estimate $d_{\rm max}$ from requiring that $h\geq h_{\rm min}$, i.e.
$d_{\rm max}=10^{2}\,\mbox{Mpc}\left(\frac{10^{-28}}{h_{\rm min}}\right)\left(\frac{m}{10^{-6}M_{\odot}}\right).$ (24)
Denoting the number density of galaxies by $n_{\rm gal}$ we can then estimate the total number of observable events from
$\dot{\mathcal{N}}_{\rm tot}\,\simeq\,\frac{4\pi}{3}d_{\rm max}^{3}n_{\rm gal}\dot{\mathcal{N}}_{\rm Gal}\,\simeq\,4\times 10^{-5}\,\mbox{yr}^{-1}\,f_{\rm PBH}\left(\frac{10^{-28}}{h_{\rm min}}\right)^{3}\left(\frac{m}{10^{-6}M_{\odot}}\right)^{2}\left(\frac{\rho_{\rm DM}^{\rm ave}}{\rho_{\rm DM}^{\rm loc}}\right)\left(\frac{N_{\rm NS}}{10^{9}}\right)\left(\frac{n_{\rm gal}}{0.01\,\mbox{Mpc}^{-3}}\right),$ (25)
where we have assumed that the Galactic event rates for the Milky Way are typical for all galaxies accounted for in $n_{\rm gal}$.
While the event rates (25) are small, even given our optimistic scaling to $h_{\rm min}=10^{-28}$ for a future gravitational wave detector, we reiterate that, close to galactic centers in particular, the dark-matter mass density $\rho_{\rm DM}$ is expected to be larger than its local value $\rho_{\rm DM}^{\rm loc}$ by several orders of magnitude.
II Stellar background models
II.1 Oppenheimer-Volkoff solutions
@@
where $R$ is the areal radius, and where we adopt geometrized units with $G=1=c$ for the remainder of this document. Here the metric coefficient $e^{2\lambda}$ is given by
@@
II.2 Transformation to isotropic coordinates
@@
In the stellar interior, however, Eq. (35) has to be integrated numerically in general, starting at the center with $r=0=R$ and fixing a constant of integration by matching to the exterior solution (36) at the stellar surface. We note that the singularity at $R=0$ on the right-hand side of (35) can be handled by adding and subtracting a term $dR/R$, i.e. by writing
@@
II.3 Constant-density star
The limit of incompressibility ($\Gamma\rightarrow\infty$) yields stars of constant density $\rho=$ const, which can be constructed analytically. For such stars we have
@@ II.3 Constant-density star
where $C=e^{\tilde{C}}$. We can now fix this constant by requiring that the isotropic radius in the interior, given by (II.3), matches its value in the exterior, given by (36), at the stellar surface $R_{*}$. This yields
@@
where $r_{*}=r(R_{*})$ is the isotropic stellar radius. We can also invert (II.3) to obtain
@@
III Primordial black holes as test particles
III.1 Geodesic equation
We track the primordial black hole’s (PBH’s) trajectory by treating it as a freely-falling test particle in the spacetime generated by a stellar background model as discussed in Section II. These spacetimes posses two Killing vectors – a time-like Killing vector $\xi_{(t)}=\partial/\partial t$ and a space-like Killing vector $\xi_{(\varphi)}=\partial/\partial\phi$ – which give rise to two constants of motion, namely the energy per unit mass
@@ III.1 Geodesic equation
Here $u^{a}$ is the PBH’s four-velocity, and the metric components can either be evaluated in terms of the areal radius $R$ (using the metric 26) or in terms of the isotropic radius $r$ (using the metric 33).
@@
(see, e.g., [54, 55] and note that the last equation is a direct consequence of 49). In (50), the time component of the four-velocity $u^{t}$ can either be found from (48), knowing $e$, or from the normalization $u_{a}u^{a}=-1$, which yields
@@
III.2 Initial conditions
We start all orbits by placing the PBH at an areal radius $R(0)=R_{\rm frac}R_{*}$ and $\varphi(0)=0$. From $R(0)$ we compute the initial isotropic radius $r(0)$ as described in Section II.2.
@@ III.2 Initial conditions
IV Perturbations of circular orbits
@@ IV.1 Stellar exterior
In this case several terms in (IV) cancel conveniently and we obtain the more familiar expression,
@@
IV.2 Stellar interior
In the stellar interior the analysis is more complicated because the metric coefficients $g_{tt}$ and $g_{RR}$ are no longer inverses of each other and hence no longer cancel each other. One approach would be to insert numerical solutions $M(R)$ and $\Phi(R)$ of the OV equations (see Section II) into (IV). Instead, we focus our analysis to a nearly homogeneous region close to the center, where we expand the density and pressure about their central values [^2] The $R^{4}$ term in the expansion of $\rho(R)$ is needed in the $M(R)\ell^{2}/R^{3}$ term of (IV) only.
@@ IV.2 Stellar interior
where $\Phi_{0}=\Phi(0)$ is a constant of integration that is fixed by matching to the exterior solution at the stellar surface. We now insert $M(R)$ and $\Phi(R)$ into (IV) and expand to order $R^{4}$ about the origin in order to identify the coefficients $p$ and $q$ in
@@
We next evaluate the derivatives $\rho_{c}^{(2)}$, $\rho_{c}^{(4)}$, and $P_{c}^{(2)}$ that we introduced in the expansions (62). In order to obtain $P_{c}^{(2)}$ we take a derivative of the stellar structure equation (30), replace the derivative of $M(R)$ using (29), and evaluate the result at $R=0$ to obtain
@@
As a consistency check we can take the Newtonian limit of (70) to find
@@
where we have used Newtonian expressions for the speed of sound $a^{2}=\Gamma K\rho^{\Gamma-1}$ as well as the Newtonian limit of (68), and where we do not distinguished between $\rho$ and $\rho_{0}$. As expected, (71) agrees with the expressions that we had found from Newtonian arguments in the main body of the paper. While it is instructive to consider the Newtonian limit, we note that the relativistic terms dominate the periastron advance for the neutron star hosts that we consider here, [^3] This can be verified by comparing the Newtonian terms with first-order post-Newtonian terms. so that a Newtonian treatment of these effects would lead to significant errors.
@@
by inserting the general relativistic expressions (70b) and (70c). Evaluating these terms requires the values ${\rho}_{c}$, $\rho_{0c}$, $P_{c}$, and $\Phi_{0}$, which depend on the stellar background model, as well as $E$ and $\ell$, which we determine for any given orbit. Results for specific stellar models and orbits are provided Table I of the main paper, where they are compared with results from integrations of the geodesic equations of Section III.
IV.3 Constant-density star
For a constant-density star, both the mass $M(R)$ and the metric coefficient $e^{\Phi}$ are known analytically, see Eqs. (39) and (41). Accordingly, we can insert these expressions into (IV) to obtain the potential $V(R)$ for all radii, without the need of an expansion about the origin. Abbreviating
@@ IV.3 Constant-density star
for all radii. Unlike the purely Newtonian prediction, this fully relativistic calculation leads to a non-vanishing periastron advance even for constant-density stars, in agreement with our numerical solutions to the geodesic equation of Section III. For the orbits listed in Table I of the main paper, for example, for which $R_{\rm frac}=0.05$, $\ell_{\rm frac}=0.99$ and $M_{*}/R_{*}=1/6$, Eq. (75) yields a periastron advance of $\Delta\phi=0.001659$.
While, for a constant-density star, an expansion about $R=0$ is not necessary, such an expansion is instructive for purposes of comparison with the approach of Section IV.2. In this case we obtain